OpenStudy (anonymous):

Use prime factorization to find the least common multiple of 120 and 252.

OpenStudy (anonymous):

120= 2*60= 2*2*30= 2*2*2*15= 2*2*2*3*5

OpenStudy (anonymous):

Now find the factorization of 252 in a similar way, go!

OpenStudy (anonymous):

2520 is lcm

OpenStudy (anonymous):

can you give me the first one please? how do you start it?

OpenStudy (anonymous):

use upside down division. much easier.

OpenStudy (anonymous):

2-120 252 2-60 126 3- 30 63 10 21 2*2*3*10*21=2520

OpenStudy (anonymous):

find prime factors of both numbers. then keep going until you can't and multiply the factors with the last pair of numbers on bottom.

OpenStudy (anonymous):

works with more than 2 numbers as well

OpenStudy (anonymous):

its 252 not 2520

OpenStudy (anonymous):

the lcm is 2520 dumbo. look it up on the internet. i'm right stupid.

OpenStudy (anonymous):

um? get off of here please?

OpenStudy (anonymous):

Ok, let's do it my way.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

your way is harder.

OpenStudy (anonymous):

nona18 im sorry im not good with this and being mean about it isnt going to help anyone.

OpenStudy (anonymous):

sorry just use my directions and by the way, prime factorization for 252 is 2 2 3 3 7

OpenStudy (anonymous):

To find the factorization of 252 start dividing by 2 as long as you can: 252= 2*126= 2*2*63 So now you can't divide by 2 anymore, and you move one to next prime: 3 2*2*63= 2*2*3*21= 2*2*3*3*7 And there's the factorization, because 2,3 and 7 are all prime.

OpenStudy (anonymous):

thank you so much thomas! you're soo helpful!

OpenStudy (anonymous):

Now to find the lcm.

OpenStudy (anonymous):

i got that part

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Al right then, you're welcome.

OpenStudy (anonymous):

i really appreciate it thomas9 thank you